OpenStudy (anonymous):

Which represent the first two terms of the sequence: a1 = -2 and an = 2(an -1)^2?

OpenStudy (anonymous):

2, -8 -8, -128 16, 1024 -2, 16

OpenStudy (ikram002p):

term \( a_1 \) is the first term term \( a_2 \) is the second term so u have \( a_1 \) try to find \( a_2 \)

OpenStudy (rsadhvika):

hint : you don't need to do any calculations for solving this problem

OpenStudy (vishweshshrimali5):

Yeah pay attention on the power of \(a_{n}\)

OpenStudy (vishweshshrimali5):

* I meant in RHS of given expression.

OpenStudy (rsadhvika):

just look at the options and use the fact that \(\large a_1 = -2\)

OpenStudy (rsadhvika):

There is only option that has the first term = \(\large -2\)

OpenStudy (ikram002p):

yeah there is a problem in options , could u plz check A @sapphirelearner1995 to make sure its 2,-8 not -2, 8 ?

OpenStudy (anonymous):

it is 2, -8 Did I write -2.8?

OpenStudy (ikram002p):

mmm idk *confused* ^_^

OpenStudy (anonymous):

lol its okay @ikram002p

OpenStudy (vishweshshrimali5):

I am going to write the question in a better way: \(\large{a_{n} = 2(a_{n-1})^2}\) This was the expression @sapphirelearner1995 right ?

OpenStudy (vishweshshrimali5):

It should be -2, 8 in place of 2,-8 in option (A).

OpenStudy (anonymous):

its\[a1 = 2\]

OpenStudy (ikram002p):

-,-

OpenStudy (vishweshshrimali5):

Ok then, also no option is right. Then, it should be 2,8 is option A.

OpenStudy (anonymous):

hehe

OpenStudy (rsadhvika):

attach the screenshot of problem

OpenStudy (vishweshshrimali5):

Your question was wrong .

OpenStudy (neer2890):

i think \[a_{n}=2(a _{(n-1)})^{2}\]

OpenStudy (vishweshshrimali5):

It was \(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 I am going to write the question in a better way: \(\large{a_{n} = \color{red}-2(a_{n-1})^2}\) This was the expression @sapphirelearner1995 right ? \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

yes @vishweshshrimali5

OpenStudy (vishweshshrimali5):

Now, see, you have this: (1) \(a_{1} = 2\) (2) \(a_{n} = -2(a_{n-1} - 1)^2\) Now, since you are given a1, so, using 2nd equation, you can find out a2 (Hint: put n = 2 in equation 2)

OpenStudy (vishweshshrimali5):

Can you find it @sapphirelearner1995 ?

OpenStudy (neer2890):

either question is wrong or the options are wrong because it doesn't suit any options.

OpenStudy (vishweshshrimali5):

@neer2890 did you check the new one : (1) \(a_{1} = 2\) (2) \(\large{a_{n} = -2(a_{n-1} - 1)^2}\)

OpenStudy (neer2890):

yup.

OpenStudy (neer2890):

a1=2 a2=-2(2-1)^2 =-2*1=-2

OpenStudy (vishweshshrimali5):

Oh no no no. Its (n-1) in subscript

OpenStudy (anonymous):

if it was like that then there wouldnt be 2 in there in first place

OpenStudy (vishweshshrimali5):

Yeah there must be given that n> 1

OpenStudy (anonymous):

\[-2(2)^{2}\] \[-2(4)=-8\]

OpenStudy (neer2890):

\[1.a _{1}=2\] \[2.a _{n}=2(a_{n-1})^{2}\]????

OpenStudy (vishweshshrimali5):

Nope thats -2 in RHS in 2nd equation

OpenStudy (anonymous):

So the answer is \[2, -8\]

OpenStudy (neer2890):

yeah ..typo :D

OpenStudy (vishweshshrimali5):

Yeah

OpenStudy (vishweshshrimali5):

Did you get it @sapphirelearner1995 ?

OpenStudy (anonymous):

K, I am going with that! Thank you everyone! I'd give ya'll metals, but I can't so if people would give others metals to, that'd be awesome!

OpenStudy (vishweshshrimali5):

No problem. We would give each other medals. :D

OpenStudy (anonymous):

I don't know if it's right (I'll tell everyone when I submit the test).

OpenStudy (vishweshshrimali5):

Ohh don't worry that's exactly correct.

OpenStudy (anonymous):

It's right :)

OpenStudy (vishweshshrimali5):

:)